OpenStudy (anonymous):

Write the quadratic function in standard form. y= -2(x+4)(x-3)

OpenStudy (marigirl):

first lets expand both the brackets.. y=-2(x^2-3x+4x-12) simplify y=-2(x^2+x-12) now multiply everything inside the bracket my -2 .. y=-2x^2-2x+24